OpenStudy (anonymous):

y1=-3sin8t-4cos8t and y2=Asin(Bt+C) transform y1 into y2

OpenStudy (sirm3d):

as in the previous problem, we expand \[A\sin( Bt + C)\] to \[A\sin(Bt) \cos C + A \sin C \cos (Bt)\]

OpenStudy (sirm3d):

and compare it to \[-3\sin(8t)-4\cos(8t)\]

OpenStudy (sirm3d):

we conclude that the coefficients of t must be the same for both expressions

OpenStudy (anonymous):

right

OpenStudy (sirm3d):

so \(B = 8\). agreed?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i wouldn't have to add them together would I be it's 2 differnet equations though would I? I just want to make sure.

OpenStudy (sirm3d):

the coefficient of \(\sin (Bt)\) in the first expression is \(A \cos C\) and the coefficient of \(\sin 8t\) is \(-3\), so we have our first relation \(-3=A\cos C\)

OpenStudy (anonymous):

i just want to make sure before we combined two to make one to satisfy both but since in this one there equal it doesn't matter right?

OpenStudy (sirm3d):

yes, we compare "similar terms"

OpenStudy (anonymous):

okay

OpenStudy (sirm3d):

doing the same with \(\cos Bt\) and \(\cos 8t\), we have our 2nd relation \[-4=A\sin C\]

OpenStudy (sirm3d):

square both sides of the equation and adding them up, we get the value of \(A\) dividing the 2nd equation by the first equation, we get an equation in \(\tan C\)

OpenStudy (sirm3d):

what did you get for \(A\)?

OpenStudy (anonymous):

I'm working it out still sorry I write stuff out step by step when I'm working a problem so I don't mess up.

OpenStudy (sirm3d):

take your time. haste makes waste.

OpenStudy (anonymous):

but I would get 25= A^2*cos^2(C)+A^2*sin^2(C)

OpenStudy (anonymous):

is this right?

OpenStudy (sirm3d):

on the right track. factor out the \(A^2\), then use a trigonometric identity.

OpenStudy (anonymous):

25=A^2(cos^2(C)+sin^2(C))

OpenStudy (sirm3d):

correct. \[\cos^2 C + \sin^2 C=1\]

OpenStudy (anonymous):

so i need to take the square root right of both sides?

OpenStudy (sirm3d):

yes, you must, to find the value of \(A\).

OpenStudy (anonymous):

A will =5

OpenStudy (anonymous):

right?

OpenStudy (sirm3d):

right. it is 5.

OpenStudy (anonymous):

I can use either equation then right to find C?

OpenStudy (sirm3d):

before you proceed with angle C have you observed that in the two problems, \[A^2 = (-4)^2+(3)^2 \quad \text {previous}\]\[A^2 = (-4)^2+(-3)^2 \quad \text {current}\]

OpenStudy (sirm3d):

these are the constants in the pair of equations \[\begin{cases} -4=A \cos C \\3 = A\sin C\end{cases}\]and \[\begin{cases} -3=A \cos C \\-4 = A\sin C\end{cases}\]

OpenStudy (anonymous):

I haven't noticed the previous one. I thought it was the current one from the beginning until you pointed it out.

OpenStudy (sirm3d):

they appear everytime in this class of trigonometric problems, so it should be steps shorter if you remember them.

OpenStudy (anonymous):

okay

OpenStudy (sirm3d):

when we're done with this problem, i'll give you an example on finding \(A\).

OpenStudy (anonymous):

okay

OpenStudy (sirm3d):

finding \(C\) is the tricky part. you have to account for the quadrant when choosing the value of C. your calculator may give you an angle but it may not be the right value for C.

OpenStudy (anonymous):

okay so c will be in the first or fourth quadrant right?

OpenStudy (sirm3d):

\(-3=5 \cos C,\quad C\) in quadrant II or III

OpenStudy (anonymous):

so i'm looking were cos would be negative then not positive right?

OpenStudy (anonymous):

that's why it's the 2 and 3?

OpenStudy (sirm3d):

\(-4=5\sin C\quad C\) in quadrant III or IV

OpenStudy (sirm3d):

yes, Q II or Q III for cosine, Q III or Q IV for sine. i have provided a sign check above.

OpenStudy (anonymous):

okay.

OpenStudy (sirm3d):